Arlene is stationary relative to the ground. Aiza is riding toward A rlene on her e-bike 5 m/s relative to the ground. Aiza throws a water balloon forward at 3 m/s relative to the e-bike. With what speed does it strike Arlene? (Solution)


1
Expert's answer
2021-11-25T10:17:16-0500

water balloon speed relative to the bike:

"v_1=v-v_2=5-v_2=3" m/s

where "v" is bike speed relative to the ground,

"v_2"  is water balloon speed relative to the ground

then water balloon strikes Arlene with speed:

"v_2=5-3=2" m/s